# Unraveling additional quantum many-body scars of the spin-$1$ XYXY model with Fock-space cages and commutant algebras

### v1

@@ IV.2.2 Reverse-engineering approach
$\begin{split}&q_{i,i{+}1}|r,n\rangle{=}0~~\forall i,~\forall n,~~~~~\text{where}\\ &q_{i,i{+}1}{=}(|0,0\rangle\langle 1,{-}1|{+}|0,0\rangle\langle{-}1,1|{+}\text{h.c.})_{i,i{+}1},\end{split}$ (23)
@@
while for r=1r{=}1 an analogous structure holds by a next-to-next nearest neighbor operator set: $q_{i,i{+}3}|1,n\rangle{=}0~~\forall i,n$. The existence of these local annihilators suffices to account for the fact that these are scars.
@@
Figure 9: Level statistics showing adjacent-gap ratio distribution $P(r_{\alpha})$ for the Hamiltonian $H_{\rm scar}{=}H_{0}{+}H_{\rm pert}$. Calculations are done for a spin-11 chain of size 2L=142L{=}14 with PBC and (J,h)=(1,1)(J,h){=}(1,1) within the symmetry-resolved sector of magnetization M=10M{=}{-}10, momentum p=0p{=}0, and inversion parity $\mathcal{I}{=}1$ (dimension=9840{=}9840). Only the bulk of the spectrum is used, excluding extensive zero modes (a) with perturbation $V_{\rm comm}^{(2)}$ [see Eq. (22)] using staggered periodic couplings $\epsilon_{i}{=}\epsilon_{i{+}2}{\in}[0,0.2]$, (b) with perturbation $V_{\rm NNNN}$ [see Eq. (26)] using $J_{3}{=}0.2$. In both cases, the distribution agrees with the analytic prediction [97] from the GOE of RMT (blue), with average $\langle r_{\alpha}\rangle{\approx}0.53$.
@@
Figure 10: Illustration of a volume-law tower state $|\mathbb{V}_{2}\rangle$ [Eq. (25)] for a chain of size 2L=82L{=}8. For ease of notation, the local basis states $|{-}1\rangle,|0\rangle,|1\rangle$ are simply denoted as $|{-}\rangle,|0\rangle,|{+}\rangle$. The state is a coherent superposition of configurations with two momentum $\pi$ bimagnons either located on distinct antipodal pairs or doubly occupying a single antipodal pair.
@@
An immediate advantage of this approach is that it naturally identifies a large class of perturbations that preserve the states $|r,n\rangle$. Introducing such perturbations is essential for realizing these states as true QMBS in a non-integrable setting, since the parent Hamiltonian $H_{0}$ has many additional symmetries, and might also be integrable in even magnetization sectors as discussed in Sec. II.1. Since any operator belonging to the algebra $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\{q_{i,i{+}1}\}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (for r>1r{>}1) or $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\{q_{i,i{+}3}\}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (for r=1r{=}1) have the states $|r,n\rangle$ as eigenstates, such operators can be used directly as admissible perturbations $V_{\rm pert}$ that preserve the scars. These range from simple nearest-neighbor sums $\sum_{i}q_{i,i+1}$ to multi-site constructions such as $\sum_{i}q_{i,i+1}q_{i+2,i+3}$ built from products of the local annihilators. Thus, the reverse engineering approach provides an alternative way to recognize that the QMBS are simultaneous eigenstates of non-commuting local operators, which in turn enables the construction of a wider class of scar-preserving perturbations. In addition, identifying local operators that annihilate these states might provide a way to understand these in the Shiraishi-Mori framework of QMBS [98], the exploration of which we leave for future work.
IV.3 Additional family: volume-entangled states
@@ IV.3.1 Structure of the states
such that $H_{XY}{=}\sum_{j}\mathcal{H}_{j}^{(L)}$. The common kernel of these operators contains not only the previously discussed $|L,n\rangle$ states but also an additional family of (unnormalized) states denoted $|\mathbb{V}_{n}\rangle$ (see Appendix C.3.2):
where $n{=}0,1,{\cdots},2L$. Note that although it may appear that these states could be generated by repeatedly applying a single operator such as $Q^{\dagger}{=}\sum_{i=1}^{L}{(-1)^{i}S^{+}_{i}S^{+}_{i+L}}$, this does not reproduce the required equal-weight structure-$(Q^{\dagger})^{n}$ produces configuration–dependent combinatorial coefficients. A short illustrative example is provided in Appendix C.3.1.
Each $|\mathbb{V}_{n}\rangle$ corresponds to a configuration with $n$ “antipodal bimagnons”, pairs of spin-raising operators acting on sites $i$ and i+Li{+}L, dispersing/rotating with momentum $\pi$ [see Fig. 10 for a schematic for $|\mathbb{V}_{n{=}2}\rangle$ for 2L=82L{=}8] on the fully polarized vacuum $|\Omega\rangle$. Note that, unlike the previously studied “bond bimagnon” scars [Eq. (9)] in Refs. [32, 33], where bimagnons are maximally local since the magnons reside on neighboring sites, here they are maximally nonlocal since the magnons are $L$ sites apart.
As these states only involve even numbers of excitations, their total magnetization is $M_{n}{=}{-}2(L{-}n)$. Thus, they are eigenstates of the full Hamiltonian $H_{0}$ with eigenvalues $E_{n}{=}{-}2h(L{-}n)$. As before, we can directly see that a simple perturbation $V_{\rm comm}^{(L)}{=}\sum_{j{=}1}^{r}\epsilon_{j}\mathcal{H}_{j}^{(r)}$ (with $\epsilon_{j}{=}\epsilon_{j{+}L}$) which annihilates all $|\mathbb{V}_{n}\rangle$. In addition, we find that the translationally invariant next-to-next-nearest-neighbor (NNNN) $XY$ Hamiltonian considered in Ref. [32]:
@@
with arbitrary coupling $J_{3}$, also preserves the entire tower $|\mathbb{V}_{n}\rangle$. Thus, the resulting total Hamiltonian $H_{\rm vol}{=}H_{0}{+}V_{\rm NNNN}$ admits $|\mathbb{V}_{n}\rangle$ as a tower of exact eigenstates, with unchanged energies $E_{n}{=}{-}2h(L{-}n)$. Moreover, the level-spacing statistics of $H_{\rm vol}$ follow the Wigner–Dyson distribution as shown in Fig. 9 (b), confirming that the model is non-integrable.
@@ IV.3.2 Entanglement properties
The entanglement properties of $|\mathbb{V}_{n}\rangle$ are markedly different from those of other scar states discussed in this work. They depend sensitively on the choice of the bipartition as illustrated in Fig. 5, where we show the bipartite EE of all $|\mathbb{V}_{n}\rangle$ states for a 2L=122L{=}12-site chain under a standard bipartition, where subsystem $A_{s}{=}\{1,2,{\cdots},L\}$ and a fine-tuned bipartition, where subsystem $A_{f}{=}\{1,L{+}1,2,L{+}2,{\cdots},L/2,L/2{+}L\}$, consisting of $L/2$ pairs of antipodal sites. Under the fine-tuned bipartition, $|\mathbb{V}_{n}\rangle$ has sub-extensive (typical scar-like) entanglement, while under the standard bipartition, $|\mathbb{V}_{n}\rangle$ (specifically mid-spectrum states of the tower) exhibits substantially higher EE than other exact scars, placing them close to the thermal continuum.
@@
$\mathcal{S}_{A_{s}}(n){=}\ln(|\mathcal{D}_{L}^{M_{n}/2}|),$ (27)
where $|\mathcal{D}_{L}^{M}|$ denotes the dimension of the spin-$1$ Hilbert space of $L$ sites with magnetization $M$ [see Eq. (5)]. Using this, we compute the EE for several representative states in the tower, specifically mid-spectrum states with $n{=}L,L{+}1,{\cdots},L{+}4$ that reside in the largest magnetization sectors $M_{n}{=}0,2,{\cdots},8$. These are the central states of the tower and exhibit the highest entanglement in the family. In Fig. 11(a), we plot the standard-cut EE of these states as a function of system size $2L$. Numerical fits reveal that the EE grows linearly with system size, consistent with a volume-law–type growth but with a logarithmic correction. In fact, one can show analytically, e.g., for the mid-spectrum state $|\mathbb{V}_{n{=}L}\rangle$ (which exhibits the maximal entanglement within the tower), EE takes the asymptotic form
$\mathcal{S}_{A_{s}}(L)\underset{\text{large }L}{\longrightarrow}L\ln(3){-}\frac{1}{2}\ln(L){+}{\rm constant}$ (28)
@@
$\begin{split}&\mathcal{S}_{A_{f}}(n){=}-\sum_{k{=}K_{\rm min}}^{K_{\rm max}}\lambda_{k}^{2}\ln\lambda_{k}^{2}\\ &{\rm where}~~~~\lambda_{k}{=}\frac{\sqrt{|\mathcal{D}_{L/2}^{k}|}\sqrt{|\mathcal{D}_{L/2}^{M_{n}/2-k}|}}{\sqrt{|\mathcal{D}_{L}^{M_{n}/2}|}}~~~{\rm and}\\ &K_{\rm min}{=}{\rm max}\left(\frac{-L}{2},\frac{M_{n}{-}L}{2}\right),~~K_{\rm max}{=}{\rm min}\left(\frac{L}{2},\frac{M_{n}{+}L}{2}\right).\end{split}$ (29)
@@
demonstrating that the fine-tuned cut obeys subextensive EE scaling at most logarithmically with system size. The numerical fits shown in Fig. 11(b) are fully consistent with this predicted logarithmic scaling.
@@ IV.3.3 Oscillations from a simple but volume-law entangled initial state
To probe the dynamical behavior associated with this special tower, we consider the coherently superposed initial state $|\psi_{\rm vol}^{\rm init}\rangle{=}\sum_{n}|\mathbb{V}_{n}\rangle$, formed by summing over all scar states in the tower. This state exhibits long-lived fidelity oscillations (see Fig. 4) under time evolution with $H_{\rm vol}$, indicating robust nonthermal dynamics. Remarkably, $|\psi_{\rm vol}^{\rm init}\rangle$ can be written exactly as a product of maximally entangled dimers between antipodal site pairs:
@@
This structure is reminiscent of the entangled antipodal pair (EAP) eigenstates discussed in Ref. [22, 23, 24], though here it appears in a distinct context: not as an eigenstate, but as an initial state that supports long-lived oscillations. This distinguishes the tower from previously studied scar families, where the initial states exhibiting oscillations are typically weakly entangled under all bipartitions. In contrast, $|\psi_{\rm vol}^{\rm init}\rangle$ is highly entangled with $\mathcal{S}_{A_{s}}{=}L\ln(3)$ under the standard bipartition, but separable for the fine-tuned antipodal bipartition, where it becomes a simple product state. This also directly explains the differences observed in the EE scaling of the constituent scar states $|\mathbb{V}_{n}\rangle$ under different bipartitions, which can be viewed as projections of $|\psi_{\rm vol}^{\rm init}\rangle$ onto fixed magnetization sectors.
Figure 12: (a) The coherently superposed state $|\psi_{\rm vol}\rangle{=}\sum_{n}|\mathbb{V}_{n}\rangle$ for 2L=122L{=}12, that exhibits long-lived fidelity oscillations. Its structure is an exact product of entangled dimers between antipodal sites, leading to volume-law entanglement under standard bipartition and vanishing entanglement under antipodal bipartition. (b) Mirror scar state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ with k=1k{=}1 for 2L=122L{=}12, formed by pairing dimers symmetrically about a reflection axis passing through sites $k$ and k+Lk{+}L. The two spins located on the symmetry axis, labeled by $m,m^{\prime}$, can take any of the nine possible two-site spin-$1$ configurations.
IV.4 Additional family: Mirror-dimer states with unconstrained spins
We now turn to a final class of exact scar eigenstates discovered through the construction of simultaneous eigenstates. These states possess a distinctive mirror-dimer structure and feature two unconstrained spins located along the reflection axis that remain completely free. To construct them, we place the spin-$1$ chain of length $2L$ on a ring and choose a reflection axis passing through an antipodal pair of lattice sites (k,k+L)(k,k{+}L). For each such axis, we define a family of mirror-symmetric operators

### v2

@@ IV.2.2 Reverse-engineering approach
$\begin{split}&q_{i,i{+}1}|r,n\rangle{=}0~~~\forall r{>}1,\forall i,~\forall n,~~~~~\text{where}\\ &q_{i,i{+}1}{=}(|0,0\rangle\langle 1,{-}1|{+}|0,0\rangle\langle{-}1,1|{+}\text{H.c.})_{i,i{+}1},\end{split}$ (23)
@@
where H.c. denotes Hermitian conjugate. For r=1r{=}1 an analogous structure holds for a next-to-next nearest neighbor operator set: $q_{i,i{+}3}|1,n\rangle{=}0~~\forall i,n$. The existence of these local annihilators suffices to account for the fact that these are scars.
@@
Figure 9: Level statistics showing adjacent-gap ratio distribution $P(r_{\alpha})$ for the Hamiltonian $H_{\rm scar}{=}H_{0}{+}H_{\rm pert}$. Calculations are done for a spin-11 chain of size 2L=142L{=}14 sites with periodic boundary conditions and (J,h)=(1,1)(J,h){=}(1,1) within the symmetry-resolved sector of magnetization M=10M{=}{-}10, momentum p=0p{=}0, and inversion parity $\mathcal{I}{=}1$ (dimension=9840{=}9840). Only the bulk (retaining eigenstates numbered, in ascending order, from 50 to 4300) of the spectrum is used, excluding extensive zero modes (a) with perturbation $V_{\rm comm}^{(2)}$ [see Eq. (22)] using staggered periodic couplings $\epsilon_{i}{=}\epsilon_{i{+}2}{\in}[0,0.2]$, (b) with perturbation $V_{\rm NNNN}$ [see Eq. (26)] using $J_{3}{=}0.2$. In both cases, the distribution agrees with the analytic prediction [92] from the Gaussian Orthogonal Ensemble of Random Matrix Theory (blue), with average $\langle r_{\alpha}\rangle{\approx}0.53$.
@@
Figure 10: Illustration of a volume-law tower state $|\mathbb{V}_{2}\rangle$ [see Eq. (25)] for a chain of size 2L=82L{=}8 sites. For ease of notation, the local basis states $|{-}1\rangle,|0\rangle,|1\rangle$ are simply denoted as $|{-}\rangle,|0\rangle,|{+}\rangle$. The state is a coherent superposition of configurations with two momentum $\pi$ bimagnons either located on distinct antipodal pairs or doubly occupying a single antipodal pair.
@@
An immediate advantage of this approach is that it naturally identifies a large class of perturbations that preserve the states $|r,n\rangle$. Introducing such perturbations is essential for realizing these states as true QMBS in a non-integrable setting, since the parent Hamiltonian $H_{0}$ has many additional symmetries. Since any operator belonging to the algebra $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\{q_{i,i{+}1}\}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (for r>1r{>}1) or $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}\{q_{i,i{+}3}\}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}$ (for r=1r{=}1) have the states $|r,n\rangle$ as eigenstates, such operators can be used directly as admissible perturbations $V_{\rm pert}$ that preserve the scars. These range from simple nearest-neighbor sums $\sum_{i}q_{i,i+1}$ to multi-site constructions such as $\sum_{i}q_{i,i+1}q_{i+2,i+3}$ built from products of the local annihilators. Thus, this reverse engineering approach provides an alternative way to recognize that the QMBS are simultaneous eigenstates of non-commuting local operators, which in turn enables the construction of a wider class of scar-preserving perturbations. In addition, identifying local operators that annihilate these states might provide a way to understand these in the Shiraishi-Mori framework of QMBS [93], the exploration of which we leave for future work.
IV.3 Additional family of scars: volume-entangled states
@@ IV.3.1 Structure of the states
such that $H_{XY}{=}\sum_{j=1}^{L}\mathcal{H}_{j}^{(L)}$. The common kernel of these operators contains not only the previously discussed $|L,n\rangle$ states but also an additional family of (unnormalized) states denoted $|\mathbb{V}_{n}\rangle$ (see Appendix C.3.2):
$|\mathbb{V}_{n}\rangle{=}\sum_{1{\leq}i_{1}{\leq}{\cdots}{\leq}i_{n}{\leq}L}({-}1)^{\sum_{k{=}1}^{n}i_{k}}\prod_{k{=}1}^{n}S_{i_{k}}^{+}S_{i_{k}{+}L}^{+}|\Omega\rangle,$ (25)
where $n{=}0,1,{\cdots},2L$. Note that although it may appear that these states could be generated by repeatedly applying a single operator such as $Q^{\dagger}{=}\sum_{i=1}^{L}{(-1)^{i}S^{+}_{i}S^{+}_{i+L}}$, this does not reproduce the required equal-weight structure of $|\mathbb{V}_{n}\rangle$-$(Q^{\dagger})^{n}$ produces configuration–dependent combinatorial coefficients. A short illustrative example to elucidate this is provided in Appendix C.3.1.
Each $|\mathbb{V}_{n}\rangle$ corresponds to a configuration with $n$ “antipodal bimagnons,” dispersing/rotating with momentum $\pi$ that is obtained by acting pairs of spin-raising operators at sites $i$ and i+Li{+}L on the fully polarized vacuum $|\Omega\rangle$ [see Fig. 10 for a schematic for $|\mathbb{V}_{n{=}2}\rangle$ for 2L=82L{=}8]. Note that, unlike the previously studied “bond bimagnon” scars [Eq. (9)] in Refs. [31, 32], where bimagnons are maximally local since the magnons reside on neighboring sites, here they are maximally nonlocal since the magnons are $L$ sites apart.
As these states only involve even numbers of excitations, their total magnetization is $M_{n}{=}{-}2(L{-}n)$. Thus, they are eigenstates of the full Hamiltonian $H_{0}$ with eigenvalues $E_{n}{=}{-}2h(L{-}n)$. As before, we can directly see that the simple perturbation $V_{\rm comm}^{(L)}{=}\sum_{j{=}1}^{L}\epsilon_{j}\mathcal{H}_{j}^{(L)}$ (with $\epsilon_{j}{=}\epsilon_{j{+}L}$) annihilates all $|\mathbb{V}_{n}\rangle$. In addition, we find that the translationally invariant next-to-next-nearest-neighbor (NNNN) $XY$ Hamiltonian considered in Ref. [31]:
@@
with arbitrary coupling $J_{3}$, also preserves the entire tower $|\mathbb{V}_{n}\rangle$. Thus, the resulting total Hamiltonian
$H_{\rm vol}{=}H_{0}{+}V_{\rm NNNN}$ (27)
admits $|\mathbb{V}_{n}\rangle$ as a tower of exact eigenstates, with unchanged energies $E_{n}{=}{-}2h(L{-}n)$. Moreover, the level-spacing statistics of $H_{\rm vol}$ follow the Wigner–Dyson distribution as shown in Fig. 9(b), confirming that the model is non-integrable.
@@ IV.3.2 Entanglement properties
The entanglement properties of $|\mathbb{V}_{n}\rangle$ are markedly different from those of other scar states discussed in this work. They depend sensitively on the choice of the bipartition as illustrated in Fig. 5, where we show the bipartite EE of all $|\mathbb{V}_{n}\rangle$ states for a 2L=122L{=}12-site chain under a standard bipartition, where subsystem $A_{s}{=}\{1,2,{\cdots},L\}$ consists of $L$ contiguous sites, and a fine-tuned antipodal bipartition, where subsystem $A_{f}{=}\{1,L{+}1,2,L{+}2,{\cdots},L/2,L/2{+}L\}$ also comprising $L$ sites consists of $L/2$ pairs of antipodal sites. Under the fine-tuned bipartition, $|\mathbb{V}_{n}\rangle$ has sub-extensive (typical scar-like) entanglement, while under the standard bipartition, $|\mathbb{V}_{n}\rangle$ (specifically mid-spectrum states of the tower) exhibits substantially higher EE than other exact scars, placing them close to the thermal continuum.
@@
$\mathcal{S}_{A_{s}}(n){=}\ln(\mathcal{D}_{L}^{M_{n}/2}),$ (28)
where $\mathcal{D}_{L}^{M}$ denotes the dimension of the spin-$1$ Hilbert space of $L$ sites with magnetization $M$ [see Eq. (5)]. Using this, we compute the EE for several representative states in the tower, specifically mid-spectrum states with $n{=}L,L{+}1,{\cdots},L{+}4$ that reside in the largest magnetization sectors $M_{n}{=}0,2,{\cdots},8$. These are the central states of the tower and exhibit the highest entanglement in the family. In Fig. 11(a), we plot the standard-cut EE of these states as a function of system size $2L$. Numerical fits reveal that the EE grows linearly with system size, consistent with a volume-law–type growth but with a logarithmic correction. In fact, one can show analytically, e.g., for the mid-spectrum state $|\mathbb{V}_{n{=}L}\rangle$ (which exhibits the maximal entanglement within the tower), EE takes the asymptotic form
$\mathcal{S}_{A_{s}}(L)\underset{\text{large }L}{\longrightarrow}L\ln(3){-}\frac{1}{2}\ln(L){+}{\rm constant.}$ (29)
@@
$\begin{split}&\mathcal{S}_{A_{f}}(n){=}-\sum_{k{=}K_{\rm min}}^{K_{\rm max}}\lambda_{k}^{2}\ln\lambda_{k}^{2},\\ &{\rm where}~~~~\lambda_{k}{=}\frac{\sqrt{\mathcal{D}_{L/2}^{k}}\sqrt{\mathcal{D}_{L/2}^{M_{n}/2-k}}}{\sqrt{\mathcal{D}_{L}^{M_{n}/2}}}~~~{\rm and}\\ &K_{\rm min}{=}{\rm max}\left(\frac{-L}{2},\frac{M_{n}{-}L}{2}\right),~~K_{\rm max}{=}{\rm min}\left(\frac{L}{2},\frac{M_{n}{+}L}{2}\right).\end{split}$ (30)
@@
demonstrating that the fine-tuned cut obeys subextensive EE scaling that grows at most logarithmically with system size. The numerical fits shown in Fig. 11(b) are fully consistent with this predicted logarithmic scaling.
Since entanglement entropy alone is not a definitive diagnostic of ETH satisfaction [i.e., if a state has volume-law entanglement entropy that does not necessarily imply that it respects ETH (but note that the contrapositive is true, i.e., if a state has area-law entanglement entropy, that does violate ETH)], the nonthermal character of these states can also be seen at the level of observable expectation values. For generic local observables, the expectation values in these states are identical to those of typical thermal states. However, for a class of structured nonlocal, but few-body observables with support only on pairs of antipodal sites such as $S^{z}_{i}S^{z}_{i+L}$, the expectation values remain anomalous, consistent with their low entanglement under fine-tuned bipartitions. This distinguishes these states from typical ETH-respecting thermal eigenstates, which usually have high entanglement across all bipartitions.
@@ IV.3.3 Oscillations from a simple but volume-law entangled initial state
To probe the dynamical behavior associated with the special tower of Eq. (25), we consider the initial state $|\psi_{\rm vol}^{\rm init}\rangle{=}\sum_{n{=}0}^{2L}|\mathbb{V}_{n}\rangle$, formed by summing over all scar states in the tower. This state exhibits long-lived fidelity oscillations (see Fig. 4) under time evolution with $H_{\rm vol}$ [defined in Eq. (27)], indicating robust nonthermal dynamics. Remarkably, $|\psi_{\rm vol}^{\rm init}\rangle$ can be written exactly as a product of maximally entangled dimers between antipodal site pairs [see Fig. 12(a)]:
@@
This structure is reminiscent of the entangled antipodal pair (EAP) eigenstates discussed in Refs. [22, 21, 23], though here it appears in a distinct context: not as an eigenstate, but as an initial state that supports long-lived oscillations. This distinguishes the tower from previously studied scar families, where the initial states exhibiting oscillations are typically weakly entangled under all bipartitions. In contrast, $|\psi_{\rm vol}^{\rm init}\rangle$ is highly entangled with $\mathcal{S}_{A_{s}}{=}L\ln(3)$ under the standard bipartition, but separable for the fine-tuned antipodal bipartition, where it becomes a simple product state. This also directly explains the differences observed in the EE scaling of the constituent scar states $|\mathbb{V}_{n}\rangle$ under different bipartitions, which can be viewed as projections of $|\psi_{\rm vol}^{\rm init}\rangle$ onto fixed magnetization sectors.
Figure 12: (a) The special initial state, $|\psi^{\rm init}_{\rm vol}\rangle{=}\sum_{n}|\mathbb{V}_{n}\rangle$, obtained by coherently superposing states in the volume-law tower $|\mathbb{V}_{n}\rangle$ defined in Eq. (25), for 2L=122L{=}12 sites, that exhibits long-lived fidelity oscillations [see Fig. 4]. Its structure is an exact product of entangled dimers between antipodal sites, leading to volume-law entanglement under standard bipartition and vanishing entanglement under antipodal bipartition. (b) Mirror scar state $|\mathbb{M}_{m,m^{\prime}}^{k}\rangle$ [see Eq. (34)] with k=1k{=}1 for 2L=122L{=}12 sites, formed by pairing dimers symmetrically about a reflection axis passing through the sites $k$ and k+Lk{+}L. The two spins located on the symmetry axis, labeled by $m,m^{\prime}$, can take any of the nine possible two-site spin-$1$ configurations.
IV.4 Additional family of scars: Mirror-dimer states with unconstrained spins
We now turn to a final class of exact scar eigenstates discovered through the construction of simultaneous eigenstates of non-commuting clustered operators. These states possess a distinctive mirror-dimer structure and feature two unconstrained spins located along the reflection axis that remain completely free. To construct them, we place the spin-$1$ chain of length $2L$ on a ring and choose a reflection axis passing through an antipodal pair of lattice sites $k$ and k+Lk{+}L. For each such axis, we define a family of mirror-symmetric operators